# DORSal: Diffusion for Object-centric

### v1

@@ 5 Experiments
We evaluate DORSal on challenging synthetic and real-world scenes in three settings: 1) we compare the ability to synthesize novel views of a scene with related approaches, 2) we analyze the capability for simple scene edits (object removal) by dropping individual object slots, and 3) we investigate the ability of DORSal to render smooth, view-consistent camera paths. Complete experimental details are available in Appendix C and additional results in Appendix D.
@@
MultiShapeNet (MSN) Sajjadi et al. (2022c) consists of scene with 16–32 ShapeNet Chang et al. (2015) objects each. The complex object arrangement, realistic rendering Greff et al. (2022), HDR backgrounds, random camera poses, and the use of fully novel objects in the test set make this dataset highly challenging. We use the version from Sajjadi et al. (2022a) (MSN-Hard). The Street View (SV) dataset contains photographs of real-world city scenes. The highly inconsistent camera pose distribution, moving objects, and changes in exposure and white balance make this dataset a good test bed for generative modeling. Street View imagery and permission for publication have been obtained from the authors Google (2007).
@@
For comparison, we focus on SRT and OSRT from the 3D scene understanding literature (Sajjadi et al., 2022a, c), and 3DiM from the diffusion literature (Watson et al., 2023). Because OSRT (Figure 1(a)) and DORSal (Figure 1(b)) leverage the same object-centric scene representation, we can compare them in terms of the quality of generated novel-views as well as the ability to perform object-level scene edits. SRT, which was previously applied to Street View and mainly differs to OSRT in terms of its architecture, does not include Object Slots as a bottleneck. We use Sup-OSRT to compute object-slots for DORSal on MultiShapeNet and plain OSRT on Street View (where ground-truth masks are unavailable).
@@ 5.1 Novel-view Synthesis
We separately train DORSal, OSRT, SRT, and 3DiM on MultiShapeNet and Street View, where DORSal and (Sup-)OSRT leverage the same set of Object Slots. We quantitatively evaluate performance at novel-view synthesis on a test set of 1000 scenes. We measure PSNR and SSIM, which capture how well each novel view matches the corresponding ground truth, though are easily exploited by blurry predictions. To address this we also measure FID (Heusel et al., 2017), which compares generated novel views to ground-truth at a distributional level, and LPIPS (VGG) (Zhang et al., 2018), which measures frame-wise similarities using deep feature embeddings.
@@ Results.
Quantitative results can be seen in Table 1 and qualitative results in Figure 4. On MultiShapeNet and Street View it can be seen how DORSal obtains slightly lower PSNR and SSIM compared to SRT and (Sup-)OSRT, but greatly outperforms these methods in terms of FID, as expected. This effect can easily be observed qualitatively in Figure 4, where SRT and OSRT render novel views that are blurry (because they average out uncertainty about the scene), while DORSal synthesizes novel-views much more precisely by ‘imagining’ some of the details, while staying close to the actual content in the scene. Compared to 3DiM, which also leverages a diffusion probabilistic model, it can be seen how DORSal yields a more favorable trade-off between FID and PSNR. Especially on Street View, where there exist large gaps between different views, 3DiM struggles as it only receives a single conditioning view during training, and primarily generates variations on its input view. We provide an additional comparison to 3DiM having access to additional GT input views at inference time, and further a comparison between DDIM and DDPM for DORSal in Appendix D.
Table 1: Novel-view synthesis. Evaluation on 1 k test scenes from MultiShapeNet and Street View.
@@
Model PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ LPIPSnormal-$\downarrow$ FIDnormal-$\downarrow$ PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ LPIPSnormal-$\downarrow$ FIDnormal-$\downarrow$
SRT 25.93 0.813 0.237 67.29 23.60 0.739 0.282 87.91
OSRT 23.35 0.719 0.330 100.7 21.19 0.614 0.410 165.1
Sup-OSRT 22.64 0.680 0.358 112.1
3DiM 18.20 0.559 0.287 10.94 12.68 0.283 0.477 15.58
DORSal 18.76 0.557 0.266 11.01 16.05 0.421 0.361 16.24
@@ 5.2 Evaluation of Object-level Edits
We evaluate the scene editing capabilities of DORSal on both MultiShapeNet and Street View and compare to (Sup-)OSRT. To remove objects from the scene and compute scene edit segmentation masks we follow the protocol described in Section 3.2. We compare the edit segmentation masks obtained in this way to the ground-truth instance segmentation mask for these scenes using ARI (Rand, 1971) and mIoU, which are standard metrics from the segmentation literature. As is common practice, we compute these metrics solely for foreground objects (indicated as FG-). Because ground-truth instance segmentations are unavailable for Street View we only report qualitative results.
@@ Results.
Table 2: Scene editing. Evaluation on 1 k test scenes of MultiShapeNet (FG-mIoU in %).
@@
OSRT (Sajjadi et al., 2022a) 43.1 79.6
Sup-OSRT (Prabhudesai et al., 2023) 50.0 75.5
DORSal (64x64) 45.8 70.0
DORSal (128x128) 35.0 64.6
Quantitative results on MultiShapeNet can be seen in Table 2 and qualitative results in Figure 5. Compared to OSRT, it can be observed how DORSal is similarly capable at object-level scene editing and attains a high degree of compositionality with regards to the Object Slots it is conditioned on. Interestingly, we find that a smaller version of DORSal training on lower-resolution views achieves significantly better scene editing performance, closely approaching the performance of the supervised OSRT baseline (which here acts as an upper limit). This suggests a cascaded model training strategy as an optimal strategy for object-level scene editing for future work, where first a 64x64 model is trained, followed by one or multiple (conditional) upsampling models to increase the visual quality while retaining the structure of the scene content (Saharia et al., 2022c; Ho et al., 2022c).
On the real-world Street View dataset, the notion of an object is much more ambiguous and, unlike for MultiShapeNet, the Object Slots provided by the OSRT encoder capture individual objects less frequently. Nonetheless, we qualitatively observe how removal of individual Object Slots in DORSal can often still result in meaningful scene edits. We show a selection of successful scene edits in Figure 6, where dropping a specific Object Slot results in the removal of, for example, a car, a street sign, a trash can, or in the alteration of a building. Not all edits, however, are meaningful and many slots have little to no effect when removed, likely because the OSRT base model often assigns multiple slots to a single object such as a car. This qualitative result, however, remains quite remarkable as the model received no instance supervision whatsoever. We provide exhaustive editing examples (incl. failure cases) in Appendix D.
Figure 5: MultiShapeNet scene edits. We remove one slot at a time in the conditioning of DORSal and render the resulting scene while keeping the initial image noise fixed. In the leftmost panel, the slot corresponding to the background is removed while all objects are present. The other panels show examples of deleted objects (highlighted in red circles) when their corresponding slot is removed.
Figure 6: Street View scene edits. Removing one slot at a time, we show the most compelling examples where objects are erased from the scene. Notably, the encircled tree in the second row is generated upon removal of a slot to fill up the now-unobserved facade previously explained by the removed slot. The ground-truth scene does not contain a tree in this position.
@@
Figure 7: Camera path rendering. Top: Example of a circular camera path rendered for DORSal (64x64) trained on MultiShapeNet. While the rendered views are mostly consistent, there can be small inconsistencies in regions of high uncertainty which result in flickering artifacts (see object highlighted in red circle). Bottom: When trained on a dataset with a mix of close-by and fully-random camera views, DORSal achieves improved consistency resulting in qualitatively smooth videos.
5.4 Ablations

### v3

@@ 5 Experiments
We evaluate DORSal on challenging synthetic and real-world scenes in three settings: 1) we compare the ability to synthesize novel views of a scene with related approaches, 2) we analyze the capability for simple scene edits: object removal and object transfer between scenes, and 3) we investigate the ability of DORSal to render smooth, view-consistent camera paths. We provide detailed ablations in Appendix C.1. Complete experimental details are available in Appendix B and additional results in Appendix C.
@@
MultiShapeNet (MSN) Sajjadi et al. (2022c) consists of scenes with 16–32 ShapeNet Chang et al. (2015) objects each. The complex object arrangement, realistic rendering Greff et al. (2022), HDR backgrounds, random camera poses, and the use of fully novel objects in the test set make this dataset highly challenging. We use the version from Sajjadi et al. (2022a) (MSN-Hard). The Street View (SV) dataset contains photographs of real-world city scenes. The highly inconsistent camera pose distribution, moving objects, and changes in exposure and white balance make this dataset a good test bed for generative modeling. Street View imagery and permission for publication have been obtained from the authors Google (2007).
@@
For comparison, we focus on SRT and OSRT from the 3D scene understanding literature (Sajjadi et al., 2022a; c), and 3DiM from the diffusion literature (Watson et al., 2023). Because OSRT (Figure 1(a)) and DORSal (Figure 1(b)) leverage the same object-centric scene representation, we can compare them in terms of the quality of generated novel-views as well as the ability to perform object-level scene edits. SRT, which was previously applied to Street View and mainly differs to OSRT in terms of its architecture, does not include Object Slots as a bottleneck. We use Sup-OSRT (Prabhudesai et al., 2023) to compute object-slots for DORSal on MultiShapeNet and plain OSRT on Street View (where ground-truth masks are unavailable).
@@ 5.1 Novel-view Synthesis
We separately train DORSal, OSRT, SRT, and 3DiM on MultiShapeNet and Street View, where DORSal and (Sup-)OSRT leverage the same set of Object Slots. We quantitatively evaluate performance at novel-view synthesis on a test set of 1000 scenes. We measure PSNR, which captures how well each novel view matches the corresponding ground truth, though is easily exploited by blurry predictions (Sajjadi et al., 2017). To address this we also measure FID (Heusel et al., 2017), which compares generated novel views to ground-truth at a distributional level, and LPIPS (VGG) (Zhang et al., 2018), which measures frame-wise similarities using deep feature embeddings.
@@ Results.
Quantitative results can be seen in Tables 1 & 2 and qualitative results in Figure 4. On MultiShapeNet and Street View it can be seen how DORSal obtains slightly lower PSNR compared to SRT and (Sup-)OSRT, but greatly outperforms these methods in terms of FID, as expected. This effect can easily be observed qualitatively in Figure 4, where SRT and OSRT render novel views that are blurry (because they average out uncertainty about the scene), while DORSal synthesizes novel-views much more precisely by ‘imagining’ some of the details, while staying close to the actual content in the scene. Notabaly, in terms of LPIPS, DORSal performs the best out of all methods that condition on object representations (and thus have the same capacity for describing the content of the scene). We compare to 3DiM, which also leverages a diffusion probabilistic model, in Table 2, where we adjust DORSal to use 256 steps of DDPM (Ho et al., 2020) sampling, similar to 3DiM. It can be seen how DORSal strictly outperforms 3DiM across all metrics. Especially on Street View, where there exist large gaps between different views, 3DiM struggles to capture the content of the target view (indicated by substantially lower PSNR and higher LPIPS) as it only receives a single conditioning view during training, and primarily generates variations on its input view. We provide an additional comparison to 3DiM having access to additional GT input views at inference time in Appendix C.
Table 1: Novel-view synthesis. Comparing DORSal to methods based on scene representations.
@@
[…]
SRT 25.93 0.237 67.29 23.60 0.282 87.91
OSRT 23.35 0.330 100.7 21.19 0.410 165.1
Sup-OSRT 22.64 0.358 112.1
DORSal 18.76 0.266 11.01 16.05 0.361 16.24
Table 2: Novel-view synthesis. Comparing DORSal to 3DiM, here both methods use DDPM.
MultiShapeNet Street View
[…]
3DiM 18.20 0.287 10.94 12.68 0.477 15.58
DORSal (DDPM) 18.99 0.265 9.00 16.36 0.356 14.62
@@ 5.2 Evaluation of Object-level Edits
We evaluate the scene editing capabilities of DORSal on both MultiShapeNet and Street View and compare to the base model, OSRT, which serves as the upper bound in our comparison. To remove objects from the scene and compute scene edit segmentation masks we follow the protocol described in Section 3.2. We compare the edit segmentation masks obtained in this way to the ground-truth instance segmentation masks for these scenes using ARI (Rand, 1971) and mIoU, which are standard metrics from the segmentation literature. As is common practice, we compute these metrics solely for foreground objects (indicated as FG-). As ground-truth instance segmentations are unavailable for Street View we only report qualitative results.
@@ Results.
Table 3: Scene editing. Evaluation on MultiShapeNet (metrics in %).
@@
OSRT 43.1 79.6
Sup-OSRT 50.0 75.5
DORSal 45.8 70.0
We find that scene editing capabilities of the base OSRT model transfer to a large degree to the object-conditioned diffusion model (DORSal), even though DORSal is not trained with object-centric architectural priors or segmentation supervision. Table 3 provides a summary of quantitative scene editing results. In our comparison Sup-OSRT (Prabhudesai et al., 2023) refers to the OSRT base model trained with segmentation supervision, which provides the object slots for DORSal, i.e. this model serves as an upper bound in terms of scene editing performance (with significantly reduced visual fidelity).
On the real-world Street View dataset, the notion of an object is much more ambiguous and, unlike for MultiShapeNet, the Object Slots provided by the OSRT encoder capture individual objects less frequently. Nonetheless, we qualitatively observe how removal of individual Object Slots in DORSal can often still result in meaningful scene edits. We show a selection of successful scene edits in Figure 6, where dropping a specific slot results in the removal of, for example, a car, a street sign, a trash can, or in the alteration of a building. We provide exhaustive editing examples (incl. failure cases) in Appendix C.
We further find that slots can be transferred between scenes, with global effects such as scene lighting and object shadows correctly modeled for transferred objects. We perform slot transfer experiments by generating a single combined scene from two separate original scenes. The combined scene is obtained by taking half of the slots (i.e. latent object representations) from Scene 1 and half of the slots of Scene 2 as conditioning information for DORSal. Consequently, DORSal produces a novel scene where some objects (incl. the background) are carried over from Scene 1, mixed with objects from Scene 2. Qualitative results are shown in Figure 5.
Figure 5: Scene editing: object transfer. We highlight several transferred objects. Note that transferred objects are rendered consistently across views (see circled objects in final row) while taking into account global illumination properties of the scene in which they are placed in (e.g. shadows are rendered correctly for transferred objects).
Figure 6: Scene editing: object removal. Removing one slot at a time, we show examples on the Street View dataset where objects are erased from the scene. Notably, the encircled tree is generated upon removal of a slot to fill up the now-unobserved facade previously explained by this slot. The original scene does not contain a tree in this position.
@@
Figure 7: Camera path rendering. Top: Example of a circular camera path rendered for DORSal (64x64) trained on MultiShapeNet. While the rendered views are mostly consistent, there can be small inconsistencies in regions of high uncertainty which result in flickering artifacts (see object highlighted in red circle). Bottom: When trained on a dataset with close-by and fully-random camera views, DORSal achieves improved consistency resulting in qualitatively smooth videos.
6 Conclusion
Acknowledgments
Ethics Statement
Appendix A Limitations
While DORSal makes significant progress, there are several limitations and open problems worth highlighting, relating to 1) lack of end-to-end training, 2) worse editing performance and consistency for high-resolution training, 3) configuration of the MultiView U-Net architecture for 3D, 4) non-local editing effects, and 5) .
As we follow the design of Video Diffusion Models (Ho et al., 2022c) for simplicity, DORSal is not end-to-end trained and is ultimately limited by the quality of the scene representation (Object Slots) provided by the separately trained upstream model (OSRT). End-to-end training comes with additional challenges (e.g. higher memory requirements), but is worth exploring in future work.
We further found that training at 128x128 resolution with our model design results in decreased editing performance compared to a 64x64 model. We also observed qualitatively worse cross-view consistency in the higher-resolution model. To overcome this limitation, one would likely have to scale the model further in terms of size (at the expense of increased compute and memory requirements) or train a cascade of models to initially predict at 64x64 resolution, followed by one or more conditional upsampling stages, as done in Video Diffusion Models (Ho et al., 2022c).
For scene editing, we find that removing individual object slots can have non-local side effects, e.g. another object or the background changing its appearance, in some cases. Furthermore, edits of individual are typically not perfect, even when trained with a supervised OSRT base model: objects are sometimes only partially removed, or removal of a slot might have no effect at all. Especially on Street View, not all edits are meaningful and many slots have little to no effect when removed, likely because the OSRT base model often assigns multiple slots to a single object such as a car. This qualitative result, however, remains remarkable as the base OSRT model received no instance supervision whatsoever.
Finally, we would like to point out how the scene editing operations that are currently supported (eg. object removal, transfer) are limited and supporting more fine-grained scene editing operations is an important direction for future work. For object-level edits, such as applying a rotation or translation, it is foreseeable how supervised co-training with language can provide an interface for this as in Michel et al. (2023). Alternatively, the object representations themselves could be disentangled to a point where information about the rotation or position of an object is isolated, and can thus be manipulated independently during generation (Biza et al., 2023; Singh et al., 2022).
Appendix B Experimental Details
B.1 Evaluation
Novel View Synthesis.  We follow the experimentation protocol outlined in Sajjadi et al. (2022a; c) and evaluate DORSal and baselines using 5 and 3 novel target views for MultiShapeNet and Street View respectively. Similarly, the OSRT base model, is trained with 5 input views on these datasets. To accommodate the U-Net architecture used in DORSal and 3DiM, we crop Street View frames to 128x128 resolution.
Evaluation of Object-level Edits.  We use a median kernel size of 7 for all edit evaluations (incl. the baselines). We evaluate models on the first 1k scenes of the MultiShapeNet dataset. For DORSal, we use an identical initial noise variable (image) for each edit to ensure consistency. We use the 64x64 DORSal architecture for this set of experiment to allow for faster sampling of all possible object edits and since we found that the lower-resolution model is less susceptible to side effects during editing (e.g. slight changes in other parts of the scene), which result in a lower edit scores for the 128x128 model (60.6 vs. 70 FG-mIoU). This suggests that a good strategy for optimal object-level control of scene content would be to train a model at 64x64 resolution followed by one or more upsampling stages Saharia et al. (2022c); Ho et al. (2022a).
Camera-path Rendering.  For camera-path rendering of many views (beyond what DORSal was trained with) we deploy the sampling procedure outlined in Section 5.3. Camera trajectories follow a circular, center-facing path starting from the first input view. Further, we generate a dataset which has a mix of both nearby views (w.r.t. previously generated views) and uniformly sampled views (at random) from the full view distribution as in MultiShapeNet. Here we train DORSal using 10 input views and 10 target views (down-sampled to 64x64 resolution) to keep a similar amount of diversity when sampling far-away as well as close-by views. 3DiM is trained similarly as for the novel-view synthesis experiments.
B.2 Model Details
B.2.1 DORSal
Conditioning.  We obtain Object Slots from a separately trained OSRT model. In the case of MultiShapeNet, we train OSRT with instance segmentation mask supervision following the approach by Prabhudesai et al. (2023): we take the alpha masks produced by the broadcast decoder to obtain soft segmentation masks, which we match using Hungarian matching with ground-truth instance masks (under an L2 objective) and finally train the model using a cross-entropy loss using the alpha mask logits on the matched target masks. For Street View, we use the default unsupervised OSRT model with a broadcast decoder, as instance masks are not available. All OSRT models use 32 Object Slots.
Slot Dropout.  Ideally, slot representations that summarize the scene should be conditionally independent given an image ${\bm{z}}_{t}$, $p({\bm{s}}_{1:K}|{\bm{z}}_{t})=\prod_{k=1,K}p({\bm{s}}_{i}|{\bm{z}}_{t})$, i.e. to be able to manipulate the presence of objects independently for editing purposes. In reality, the slot representations may respect this assumption to varying degrees, with an OSRT model trained with instance-level supervision (Sup-OSRT) being more likely to achieve this. However, even if slots would exclusively bind to particular regions of the encoded input views that correspond to individual objects, slots may still share information as the input view encoder has a global receptive field. To mitigate this issue, we experimented with dropping slots from the conditioning set independently following a Bernoulli rate set as a hyper-parameter $\lambda_{sd}$. In this case the model sees slot subsets at training time (such that edits are now effectively in-distribution). While we found that this slightly affected Edit FG-ARI results for MultiShapeNet in a negative way, we found that it qualitatively resulted in more consistent object edits on Street View. See Appendix C.1 for a comparison. Unless otherwise mentioned we report results using $\lambda_{sd}=0$ for MultiShapeNet and $\lambda_{sd}=0.2$ for Street View.
Network Architecture.  For the DORSal we follow the architecture of Ho et al. (2022a), which is a U-Net that has axial attention over time, whose specification is as follows:
  • The inputs are the noisy set of $L$ target views (acting as the noisy video). The inputs are processed at multiple attention resolutions, each corresponding to a “level”, followed by a spatial down-sampling by a factor of two. Each level in the downward path is composed of three ResNet blocks having the amount of channels as indicated in Table 4. The middle stage consists of a single ResNet block (keeping the number of channels constant). The upward path mimics the downward path in reverse and has a residual connection to the corresponding block in the downward path. Attention takes place only at the third level (spatial resolution 16x16 after each of the ResNet blocks) in the downsample, middle and upsample paths, using a head dimensionality of 64 and 128 for input resolution 64x64 and 128x128 respectively.
  • The UNet is further conditioned with embeddings of the slots, target pose and diffusion noise level. The individual object slots are projected and broadcasted across views, where they are combined with the target camera pose (after projection) for each of the L views. We use sinusoidal pose embeddings of absolute camera rays, identical to the setup in OSRT (Sajjadi et al., 2022a). We apply this conditioning in the same way that text is integrated into the U-net: we attention-pool the conditioning embeddings into a single embedding and combine it with the embedding for the diffusion noise level for modulating U-Net feature maps via FiLM (Perez et al., 2018). Further, we use cross-attention as indicated above to attend to the conditioning embeddings derived from the object slots and camera poses.
A key difference is that DORSal does not require text conditioning cross attention layers, and it solely uses slot embeddings augmented with camera poses as indicated above. Further, notice how the architecture sizes we use are small compared to Ho et al. (2022a) as can be seen in Table 4. The U-Net on resolutions of 128 $\times$ 128 uses patching to avoid memory expensive feature maps. For the 16 ×$\times 16$ resolution, the ResBlocks use per-view self-attention and between-views cross-attention.
For details about the OSRT encoder used to compute the frozen object representations, we refer to Section B.2.3 below.
Table 4: DORSal U-Net architecture details.
U-Net 64 192, 384, 576 3 $16\times 16$ No
U-Net 128 256, 512, 1024 3 $16\times 16$ $2\times 2$
Training.  We adopt a similar training set-up to Ho et al. (2022c), using a cosine-shaped noise schedule, a learning rate with a peak value of 0.00003 using linear warm-up for 5000 steps, optimization using Adam with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and EMA decay for the model parameters. We train with a global batch size of 8 and classifier-free guidance with a conditioning dropout probability of $0.1$ (and an inference guidance weight of $2$). We report results after training for 1 000 000 steps. For MultiShapeNet, we use Object Slots from Sup-OSRT (i.e. supervised) and do not use slot dropout during training. For Street View, we use Object Slots from OSRT (i.e. unsupervised) and use a slot dropout probability of 0.2, which we found to improve editing quality on this dataset (compared to no slot dropout).
Camera-Path Sampling.  We leverage the iterative nature of the generative denoising process to create smooth transitions as well as global consistency between frames. Our technique is summarized in Figure 8: we iterate over a total of 25 stages (i.e. 8 steps per stage when using 200 denoising steps) and we interleave 3 types of shuffling (1 type per stage) to achieve both local and global consistency. The types of shuffling (as described in Section 3.2) are as follows: 1) no shuffle (identity), to allow the model to make blocks of the context length consistent; 2) shift the frames in time by about half of the context length, which puts frames together with new neighbours in their context, allowing the model to create smooth transitions; 3) shuffle all frames with a random permutation, to allow the model to resolve inconsistencies globally.
B.2.2 3D Diffusion Model (3DiM)
We compare to 3DiM, which is a pose-conditional image-to-image diffusion model for generating novel views of the same scene (Watson et al., 2023). During training, 3DiM takes as input a pair of views of a static scene (including their poses), where one of the views (designated as the “target view”) is corrupted with noise. The training objective is to predict the Gaussian noise that was used to corrupt the target view. During inference, 3DiM makes use of stochastic conditioning to generate 3D-consistent views of a scene. In particular, given a small set of $k$ conditioning views and their camera poses (typically $k=1$), a new view for a given target camera pose is generated by conditioning on a randomly selected view from the conditioning set at each denoising step. Each time a new view is generated, it is added to the conditioning set. For additional details, including code, we refer to Sections 6 & 7 in Watson et al. (2023).
[…]
Training.  We adopt the same training set-up as in the 3DiM paper, which consist of a cosine-shaped noise schedule, a learning rate with peak value of 0.0001 using linear warm-up for 10M samples, optimization using Adam with $\beta_{1}=0.9$ and $\beta_{2}=0.99$, and EMA decay for the model parameters. We train with a batch size of 128 and classifier-free guidance $10\%$ with a weight of $3$, as was done for the experiment on SRN cars in their paper. We report results after training for 320 000 steps.
Sampling.  We generate samples in the same way as in the 3DiM paper, using 256 DDPM denoising steps and clip to [-1, 1] after each step.
B.2.3 SRT & OSRT
SRT was originally proposed by Sajjadi et al. (2022c) with Set-Latent Scene Representations (SLSR) and subsequently adapted to Object Slots for OSRT Sajjadi et al. (2022a). At the same time, a few tweaks were made to the model, e.g. by using a smaller patch size and a larger render MLP Sajjadi et al. (2022a). For all our experiments (SRT and OSRT), we use the improved architecture from the OSRT paper. We reproduce several key details for the encoder, which is used to compute the object representations for DORSal, and refer to Appendix A.4 in Sajjadi et al. (2022a) for additional model and training details.
  • The encoder consists of a CNN with 3 blocks, each with 2 convolutional layers and ReLU activations. The first convolution in each block has stride 1, the second has stride 2. It begins with 96 channels, which are doubled with every strided convolution. The final activations are mapped with a 1x1 convolution (i.e. a per-patch linear layer) to 768 channels.
  • The CNN is followed by an encoder transformer, using 5 pre-normalization layers with self-attention (Xiong et al., 2020). Each layer has hidden size 768 (12 heads, each with 64 channels), and the MLPs have 1 hidden layer with 1536 channels and GELU activations (Hendrycks & Gimpel, 2016).
  • The encoder transformer is followed by a Slot Attention module (Locatello et al., 2020) using 1536 dimensions for slots and embeddings in the attention layers. The MLP doubles the feature size in the hidden layer to 3072. We use a single iteration of Slot Attention with 32 slots.
We use identical encoder architectures between OSRT and DORSal. Following Sajjadi et al. (2022c), we seperately train SRT and OSRT for $\sim$4M steps for each dataset.
B.3 Compute and Data Licenses
Appendix C Additional Results
C.1 Ablations